OpenStudy (anonymous):

ernie has an average of 88 on the five tests he has taken so far this semester what is the minimum grade ernie must earn on the next test to earn an average grade of 90

OpenStudy (texaschic101):

(88 + x) / 2 = 90 88 + x = 180 x = 180 - 88 x = 92 I think I did that right :)